OpenStudy (anonymous):

Do you know how to do a problem such as 4^h-k/4h+k?

OpenStudy (anonymous):

no . (4^h-k)/(4^h+k)

OpenStudy (anonymous):

\[\frac{4^{h-k}}{4^{h+k}}\] ?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

\[\large \frac{4^{h-k}}{4^{h+k}}\] \[\large 4^{(h-k)-(h+k)}\] \[\large 4^{h-k-h-k}\] \[\large 4^{-2k}\] \[\large \frac{1}{4^{2k}}\] \[\large \frac{1}{(4^2)^k}\] \[\large \frac{1}{16^k}\] So \[\large \frac{4^{h-k}}{4^{h+k}}=\frac{1}{16^k}\]

OpenStudy (anonymous):

thank you so much